OpenStudy (anonymous):

Given that (2,1) is one of the solution of simultaneous equations 3x+py=4 and x^2+ky+2x=11 where p and k are constants.Find the value of k,of p and the other solution.

OpenStudy (anonymous):

I found p=-2 and k=3

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

(2,1) is one of the solution but i don't know how to find the other solution

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

@Kainui

OpenStudy (ribhu):

value of p can be obtained by substituting x,y as 2,1 in the linear equation.

OpenStudy (ribhu):

and k can be obtained by putting x=2 in the quadratic equation, then the quadratic can be solved so other root will also be known.

OpenStudy (anonymous):

ya,i found already for value of p and k

OpenStudy (anonymous):

p=-2 and k=3

OpenStudy (anonymous):

but i don't know how to find the other solution

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

p = -2 k = 3 are right! good job :) lets find the other solution

OpenStudy (anonymous):

should we insert p=-2 and k=3 into the equation?

ganeshie8 (ganeshie8):

thats a good idea, so the given equations ` 3x+py=4 and x^2+ky+2x=11 ` become : \[3x-2y = 4\\x^2+3y+2x=11\]

OpenStudy (anonymous):

\[x=\frac{ 2y+4 }{ 3 }\]

ganeshie8 (ganeshie8):

may be isolate \(y\) instead

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

so that substitution becomes simple

OpenStudy (anonymous):

\[-2y=4-3x\]\[y=-\frac{ 4-3x }{ 2 }\]\[y=\frac{ 3x-4 }{ 2 }\]

ganeshie8 (ganeshie8):

looks good, plug that in second equation

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[x^2+3(\frac{ 3x-4 }{ 2 })+2x=11\]

OpenStudy (anonymous):

\[x^2+\frac{ 9x-12 }{ 2 }+2x=11\]

OpenStudy (anonymous):

\[\frac{ 2x^2 }{ 2 }+\frac{ 9x-12 }{ 2 }+\frac{ 4x }{ 2 }=11\]

OpenStudy (anonymous):

\[2x^2+13x-12=22\]

OpenStudy (anonymous):

\[2x^2+13x-34=0\]

ganeshie8 (ganeshie8):

do u remember this ? product of roots = c/a ?

OpenStudy (anonymous):

-34/2

OpenStudy (anonymous):

-17

ganeshie8 (ganeshie8):

yes that product of roots has to eqal -34/2

ganeshie8 (ganeshie8):

\[\alpha \beta = -17\]

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

and we know that one root is \(2\)

ganeshie8 (ganeshie8):

so : \[2 \beta = -17\]

ganeshie8 (ganeshie8):

\[ \beta = -\dfrac{17}{2}\]

ganeshie8 (ganeshie8):

so the x value of solution is \(\large x = -\dfrac{17}{2}\)

ganeshie8 (ganeshie8):

plug this in first equation and solve \(y\)

OpenStudy (anonymous):

\[y=\frac{ 3x-4 }{ 2 }\]\[y=\frac{ 3(-\frac{ 17 }{ 2 })-4 }{ 2 }\]\[y=-\frac{ 59 }{ 4 }\]

OpenStudy (anonymous):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

yw:)

ganeshie8 (ganeshie8):

that looks perfect!

OpenStudy (anonymous):

Thnx :)